# Structured World Modeling via Semantic Vector Quantization

### v1

@@ 5.2.2 CLEVR-Hard Property Comparison
6 Conclusion and Limitations
In this work, we introduced the Semantic Vector-Quantized Variational Autoencoder. Unlike traditional vector quantization methods, our model can obtain semantic neural discrete representations, capturing the rich structure of the objects in a scene. We showed that by training a prior over these semantic discrete tokens, we are able to generate multi-object scenes that follow the underlying data distribution. These semantic discrete representations are also useful for downstream tasks, outperforming the representations from patch-based discretation methods. While our model is only evaluated on static images, an interesting future direction would be to apply our method to videos to predict future frames. This may be fruitful for modeling longer video sequences since SVQ can compress each frame into a set of latents that only depend on the number of objects in the scene.
Limitations. While our method can learn semantic discrete representations and is capable of using these representations to generate images of higher visual fidelity than previous object-centric methods such as GENESIS (Engelcke et al., 2020; 2022), it is still only shown to work well on synthetic datasets with similar visual complexity as previous work (Singh et al., 2023). Although scaling unsupervised object-centric models to more realistic datasets is not a focus of this work, further improving our model so that it can work well on more realistic scenes is an important avenue of future research. Another limitation of our model is that our latent representations are all discrete. Although our visual world does consist of many discrete concepts, factors such as position and pose are continuous. It would be interesting to explore ways to combine continuous and discrete factors to better model realistic scenes.
7 Ethics Statement
@@
8 Reproducibility Statement
In addition to details about our model described in Section 3.1, we provide additional implementation details in Appendix B, including detailed hyperparameters used in our experiments. We will also release the source code upon acceptance of the paper.
9 Acknowledgments
@@
Appendix A Additional Experimental Results
@@ A.1 Codebook Analysis
A.2 Comparison with Slot-Level Quantization
@@
A.3 Prior Model Capacity for dVAE
In order to evaluate whether or not a larger capacity prior may improve the results for the patch-based dVAE baseline, we ran ablations using larger transformers for the dVAE prior on the CLEVR-Hard dataset. The results are presented in Table 6. We see that while increasing the size of the transformer for the prior does slightly improve the FID for dVAE, it still underperforms when compared to SVQ, indicating that simply scaling the dVAE prior may not be sufficient to match SVQ performance.
@@ A.3 Prior Model Capacity for dVAE
SVQ (8-layer) 43.12
A.4 Number of Blocks Ablation
@@
A.5 Codebook Size Ablation
@@
A.6 FG-ARI Segmentation Results
Table 9 shows the Foreground Adjusted-Rand-Index (FG-ARI) metric for the CLEVR datasets for different slot-based models. We see that when compared to SysBinder, SVQ performs similarly in terms of FG-ARI on CLEVR-Easy and CLEVR-Hard and slightly underperforms on CLEVR-Tex. Compared to vanilla Slot Attention, SVQ achieves higher FG-ARI on all 3 datasets.
@@ A.6 FG-ARI Segmentation Results
Slot Attention SLATE SysBinder SVQ
@@
A.7 Experiments on Google Scanned Objects
In order to evaluate SVQ on a more realistic dataset, we use a dataset where the objects are taken from the Google Scanned Objects (Downs et al., 2022). Specifically, we use the objects from the "Shoe" and "Bottles and Cans and Cups" categories. We evaluate on two versions of SVQ: SVQ-small uses the same hyperparameters as we used for the CLEVR-Hard dataset (see Table 11) and SVQ-large increases the codebook size to 256 and increases the size of the transformer decoder from 8 layers, 4 heads, model size 192 to 16 layers, 8 heads, model size 512.
We present the FID results in Table 10 and qualitative samples in Figure 10. We see that while SVQ-small is able to generate objects from the dataset, the objects are smoothed out, resulting in a high FID score. Increasing the model size to SVQ-large significantly improves the quality of the generated scenes and the FID score, providing some evidence that SVQ can be scaled to work on more realistic datasets.
@@ A.7 Experiments on Google Scanned Objects
SVQ-small 114.16
SVQ-large 72.68
Figure 10: Samples for SVQ on the Google Scanned Objects dataset. Scaling to a larger model size noticeably improves the quality of the samples.
Appendix B Implementation Details
B.1 Training and Implementation Details.
We use input images of 64x64 resolution for the 2D Sprites datasets and 128x128 for the CLEVR datasets. Each model is trained on NVIDIA Quadro RTX 8000 GPUs with 48GB memory and we use half-precision floating-point format. We train SVQ for 400k iterations which takes around 80 hours for the CLEVR datasets and 50 hours for the 2D datasets. We then train the prior for 1 million iterations which takes around 40 hours. For the 2D Sprites dataset, similar to (Yoon et al., 2023), we first train the underlying models on a dataset of random shapes without any relationship between the objects. We then train the prior models on the odd-one-out datasets.
B.2 Hyperparameters
@@
B.3 Prior Models
For the SVQ and DVAE prior models, we use a transformer architecture with 8 layers, 4 heads, model dimension 192, feedforward dimension 768, and a dropout probability of 0.1. We use a learning rate of 0.0003 and 30,000 warmup steps. For VQ-VAE, we use a 20-layers PixelCNN prior, as proposed in the original paper (van den Oord et al., 2017).
B.4 Downstream Models

### v2

@@ 5.2.2 CLEVR-Hard Property Comparison
6 Conclusion
In this paper, we introduce the Neural Language of Thought Model (NLoTM). This is the first model that satisfies our proposed desiderata of a neural network model of the LoTH: (1) Compositional semantic representation, (2) Symbol-like discrete concept abstraction, (3) Probabilistic neural grammar to generate samples compositionally. The NLoTM consists of two main components: Semantic Vector Quantized Variational Autoencoder, which learns hierarchical and composable discrete representations aligned with objects and their properties; and Autoregressive Language of Thought Prior, which learns to generate semantic concept tokens compositionally in a probabilistic manner. We evaluated NLoTM on several 2D and 3D image datasets. It demonstrated superior performance in downstream tasks, out-of-distribution generalization, and image generation quality compared to patch-based VQ-VAE and continuous object-centric representations. By further developing and refining models like NLoTM, we believe that we can continue to make progress in creating AI systems that exhibit more human-like understanding and generalization capabilities.
Ethics Statement
@@ 6 Conclusion
Reproducibility Statement
In addition to details about our model described in Section 3.1, we provide additional implementation details in Appendix C, including detailed hyperparameters used in our experiments. We will also release the source code upon acceptance of the paper.
Acknowledgments
@@
Appendix A Limitations
While our method can learn semantic discrete representations and is capable of using these representations to generate images of higher visual fidelity than previous object-centric methods such as GENESIS (Engelcke et al., 2020; 2022), it is still only shown to work well on synthetic datasets with similar visual complexity as previous work (Singh et al., 2023). Although scaling unsupervised object-centric models to more realistic datasets is not a focus of this work, further improving our model so that it can work well on more realistic scenes is an important avenue of future research. Another limitation of our model is that our latent representations are all discrete. Although our visual world does consist of many discrete concepts, factors such as position and pose are continuous. It would be interesting to explore ways to combine continuous and discrete factors to better model realistic scenes.
Appendix B Additional Experimental Results
@@ B.1 Codebook Analysis
B.2 Comparison with Slot-Level Quantization
@@
B.3 Prior Model Capacity for dVAE
In order to evaluate whether or not a larger capacity prior may improve the results for the patch-based dVAE baseline, we ran ablations using larger transformers for the dVAE prior on the CLEVR-Hard dataset. The results are presented in Table 6. We see that while increasing the size of the transformer for the prior does slightly improve the FID for dVAE, it still underperforms when compared to NLoTM, indicating that simply scaling the dVAE prior may not be sufficient to match NLoTM performance.
@@ B.3 Prior Model Capacity for dVAE
NLoTM (8-layer) 43.12
B.4 Number of Blocks Ablation
@@
B.5 Codebook Size Ablation
@@
B.6 FG-ARI Segmentation Results
Table 9 shows the Foreground Adjusted-Rand-Index (FG-ARI) metric for the CLEVR datasets for different slot-based models. We see that when compared to SysBinder, NLoTM performs similarly in terms of FG-ARI on CLEVR-Easy and CLEVR-Hard and slightly underperforms on CLEVR-Tex. Compared to vanilla Slot Attention, NLoTM achieves higher FG-ARI on all 3 datasets.
@@ B.6 FG-ARI Segmentation Results
Slot Attention SLATE SysBinder NLoTM
@@
B.7 Experiments on Google Scanned Objects
In order to evaluate NLoTM on a more realistic dataset, we use a dataset where the objects are taken from the Google Scanned Objects (Downs et al., 2022). Specifically, we use the objects from the "Shoe" and "Bottles and Cans and Cups" categories. We evaluate on two versions of NLoTM: NLoTM-small uses the same hyperparameters as we used for the CLEVR-Hard dataset (see Table 11) and NLoTM-large increases the codebook size to 256 and increases the size of ALP model from 8 layers, 4 heads, model size 192 to 16 layers, 8 heads, model size 512.
We present the FID results in Table 10 and qualitative samples in Figure 10. We see that while NLoTM-small is able to generate objects from the dataset, the objects are smoothed out, resulting in a high FID score. Increasing the model size to NLoTM-large significantly improves the quality of the generated scenes and the FID score, providing some evidence that NLoTM can be scaled to work on more realistic datasets.
@@ B.7 Experiments on Google Scanned Objects
NLoTM-small 114.16
NLoTM-large 72.68
Figure 10: Samples for NLoTM on the Google Scanned Objects dataset. Scaling to a larger model size noticeably improves the quality of the samples.
Appendix C Implementation Details
C.1 Training and Implementation Details.
We use input images of 64x64 resolution for the 2D Sprites datasets and 128x128 for the CLEVR datasets. Each model is trained on NVIDIA Quadro RTX 8000 GPUs with 48GB memory and we use half-precision floating-point format. We train SVQ for 400k iterations which takes around 80 hours for the CLEVR datasets and 50 hours for the 2D datasets. We then train the ALP prior for 1 million iterations which takes around 40 hours. For the 2D Sprites dataset, similar to (Yoon et al., 2023), we first train the underlying models on a dataset of random shapes without any relationship between the objects. We then train the prior models on the odd-one-out datasets.
C.2 Hyperparameters
@@
C.3 Prior Models
For the DVAE and ALP prior models, we use a transformer architecture with 8 layers, 4 heads, model dimension 192, feedforward dimension 768, and a dropout probability of 0.1. We use a learning rate of 0.0003 and 30,000 warmup steps. For VQ-VAE, we use a 20-layers PixelCNN prior, as proposed in the original paper (van den Oord et al., 2017).
C.4 Downstream Models